• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 16:15:24 on May31,2002

    Entry number 4300

    keyword=shift Summary

    Shift Summary

    0800: Run 1034 starting. KIN Opt_03 - Single foil target optics run
    0845: Run 1034 ended, 335k events. MCC needs some time to tune beam
    MCC says it may take up to few hours to check out beam


    14:45 MCC says beam may come in the next hour
    Enter into hall, Seive position checked - right on (Mark Stevens).
    Removed camera, checked gas pressure valves for cerenkov (Reitz)

    End of Shift -- Still no beam